EULER SUMS AND CONTOUR INTEGRAL REPRESENTATIONS

PHILIPPE FLAJOLET AND BRUNO SALVY

Abstract. This paper develops an approach to the evaluation of Euler sums that involve harmonic numbers, either linearly or nonlinearly. We give explicit formulæ for several classes of Euler sums in terms of Riemann zeta values. The approach is based on simple contour integral representations and residue computations.